OpenStudy (anonymous):

write the slope intercept equation for the line with the given slope that contains the given slope. m= 8/3 ; (3,-4)

OpenStudy (anonymous):

we have: y = (8 / 3) x + b plug in (3, -4) to the equation, we get (-4) = 8 / 3 * 3 + b solving we get b = -12 solution: y = (8 / 3) x - 12

OpenStudy (anonymous):

could you show me how you solved for -12. sorry i get very confused with math

OpenStudy (anonymous):

general equation is y-y1 = m(x-x1) where (x1,y1) is the given point and m is the given slope

OpenStudy (anonymous):

after plugging in the points (3, -4), we have (-4) = (8 / 3) * 3 + b proceed by adding [- (8 / 3) * 3] on both side of the equation on the left side we have: (-4) - (8 / 3) * 3 = -12 on the right side we have (8 / 3) * 3 - (8 / 3) * 3 + b = b thus -12 = b

OpenStudy (anonymous):

thank you much!

OpenStudy (anonymous):

you're welcome